Phelm Corporation sold a piece of equipment for cash and recognized a gain of $5,000. The original cost was $34,000 and the accumulated depreciation just prior to the sale totaled $19,000. What amount will appear in the investing activities section of the statement of cash flows?

$20,000
$54,000
$10,000
$15,000
Answer #1

Book value of equipment = Cost-Accumulated depreciation = 34000-19000 = 15000

Sale value of equipment = Book value of equipment+gain on equipment = 15000+5000 = $20000

So answer is a) $20000